40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1685     Entry time: Tue Jul 26 11:33:49 2016
Author: awade and Antonio 
Type: Notes 
Category: North Cavity 
Subject: North cavity temperature dropped 

The north cavity temperature was dropped this morning. The voltage of the heater power suply was dropped from 10.7 V to 10.4 V.

ELOG V3.1.3-